A SURVEY OF 100 RESPONDENTS WAS CONDUCTED TO STUDY PROPORTION PEOPLE SNORING IN THEIR SLEEP. Approximately 40% people claim that they do. Using this information, construct 90% confidence interval for the proportion of people that do snore in their sleep.


1
Expert's answer
2021-02-11T13:46:45-0500

"n=100, p=0.4, q=1-0.4=0.6"

"x=0.4\cdot100=40"

Confidence level:

"CL=0.90, \alpha=1-0.9=0.1, \alpha/2=0.05, Z_{0.05}=1.645"

The error bound for a proportion:

"EBP=Z_{\alpha/2}\sqrt{\frac{pq}{n}}=1.645\sqrt{\frac{0.4\cdot0.6}{100}}=0.081"

"p-EBP=0.4-0.081=0.319"

"p+EBP=0.4+0.081=4.081"

The confidence interval for the true binomial population proportion:

"(0.319,4.081)"

We estimate with 90% confidence that the true percent of people that do snore in their sleep is between 31.9% and 40.81%.